# Forcing Planets to Evolve: The Relationship Between Uranus and Neptune at Late Stages of Dynamical Evolution

### v1

@@ Appendix A Modifications to the equations of motion
Modifications to Mercury 6.2 in the way we are using them, were done by Dawson & Murray-Clay (2012) and Wolff et al. (2012) and are described in the Appendix of Wolff et al. (2012). We reiterate those modifications in this section with specifics and discuss additional modifications to the equations of motions to allow for artificial evolution of the phase angles argument of pericenter and longitude of ascending node. Modifications to REBOUND are explained below as well. We note that these modifications are non-symplectic, even for a symplectic integrator, thus, the modifications must represent small perturbations on the planet for this to be valid. These modifications were tested for the MVS and HYBRID integrators within Mercury and for the WHFAST and MERCURIUS integrators within REBOUND; the tests were run for systems with a star and 1, 2, or 4 planets.
@@
    Then, we tranform the coordinates back to democratic heliocentric with reb_integrator_mercurius_inertial_to_dh(r);
  • Lastly, with the addition of the new attributes to the particles struct, the simulation archive needed significant changes.
@@
To derive the additional terms, we allow all orbital elements to vary with time and define them to be independent from each other. The position vector components ($x$,$y$,$z$) are given in equation 3, as a function of the orbital elements. The variable $r$ is the distance of the planet from the central body, $\Omega$ is the longitude of ascending node in the xy plane, $\omega$ is the arguments of periapse, $f$ is the true anamoly, and $i$ is the inclination.
$x=r\cos{\Omega}\cos{(\omega+f)}-r\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (15)
$y=r\sin{\Omega}\cos{(\omega+f)}+r\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (16)
$z=r\sin{i}\sin{(\omega+f)}$ (17)
The velocity vetctor componenets are given by ($\dot{x}$, $\dot{y}$, and $\dot{z}$)
$\dot{x}$ $=\dot{r}\cos{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (18)
$-(r\dot{\omega}+r\dot{f})\cos{i}\sin{\Omega}\cos{(\omega+f)}$
$\dot{y}$ $=\dot{r}\sin{\Omega}\cos{(\omega+f)}+r\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (19)
$-r\sin{i}\dot{I}\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\sin{\Omega}\dot{\Omega}\sin{(\omega+f)}$
$+(r\dot{\omega}+r\dot{f})\cos{i}\cos{\Omega}\cos{(\omega+f)}$
$\dot{z}$ $=\dot{r}\sin{i}\sin{(\omega+f)}+r\cos{i}\dot{I}\sin{(\omega+f)}$ (20)
$+(r\dot{w}+r\dot{f})\sin{i}\cos{(\omega+f)}$
The ficticious x-component "velocity" and "acceleration" used to update the planets’ position and velocity in MERCURY are given by:
$\large\dot{x}=\frac{dx}{dt}=\frac{\partial x}{\partial a}\dot{a}+\frac{\partial x}{\partial e}\dot{e}+\frac{\partial x}{\partial i}\dot{i}+\frac{\partial x}{\partial\omega}\dot{\omega}+\frac{\partial x}{\partial\Omega}\dot{\Omega}+\frac{\partial x}{\partial f}\dot{f}$ (21)
$\large\ddot{x}=\frac{d\dot{x}}{dt}=\frac{\partial\dot{x}}{\partial a}\dot{a}+\frac{\partial\dot{x}}{\partial e}\dot{e}+\frac{\partial\dot{x}}{\partial i}\dot{i}+\frac{\partial\dot{x}}{\partial\omega}\dot{\omega}+\frac{\partial\dot{x}}{\partial\Omega}\dot{\Omega}+\frac{\partial\dot{x}}{\partial f}\dot{f}$ (22)
where similar expressions are derived for the other coordinates. These equations do not describe the entire evolution of the planets; they are additional terms due to their orbital element evolution. In MERCURY, these are added onto the existing acceleration and velocity terms in the code. Note that we do not include the $\dot{f}$ term since that is inherent to the Keplerian motion of the planet, and we do not update that value. These partial derivatives include $r$, $\dot{r}$, $r\dot{f}$ and their derivatives w.r.t. $a$,$e$, $f$:
$r=\frac{a(1-e^{2})}{1+e\cos{f}}$ (23)
$r\dot{f}=\sqrt{\frac{M}{a}}\frac{1+e\cos{f}}{\sqrt{1-e^{2}}}$ (24)
$\dot{r}=\frac{\partial r}{\partial a}\dot{a}+\frac{\partial r}{\partial e}\dot{e}+\frac{\partial r}{\partial f}\dot{f}$ (25)
$\dot{f}=\frac{r\dot{f}}{r}=\sqrt{\frac{M}{a^{3}}}\frac{(1+e\cos{f})^{2}}{(1-e^{2})^{3/2}}$ (26)
$\frac{\partial r}{\partial a}=\frac{1-e^{2}}{1+e\cos{f}}=\frac{r}{a}$ (27)
$\frac{\partial r}{\partial e}=-\frac{2ae}{1+e\cos{f}}-\frac{a(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}$ (28)
$\frac{\partial r}{\partial f}=\frac{a(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}$ (29)
$\frac{\partial r\dot{f}}{\partial a}=\frac{-r\dot{f}}{2a}=-\frac{1}{2}\sqrt{\frac{M}{a^{3}}}\frac{1+e\cos{f}}{\sqrt{1-e^{2}}}$ (30)
$\frac{\partial r\dot{f}}{\partial e}=\frac{r\dot{f}(e+\cos{f})}{(1-e^{2})(1+e\cos{f})}=\sqrt{\frac{M}{a}}\frac{e+e\cos{f}}{(1-e^{2})^{3/2}}$ (31)
$\frac{\partial r\dot{f}}{\partial f}=\frac{-a\sqrt{M/a^{3}}e\sin{f}}{\sqrt{1-e^{2}}}$ (32)
$\frac{\partial\dot{f}}{\partial a}=-3/2\frac{\sqrt{M/a^{5}}(1+e\cos{f})^{2}}{(1-e^{2})^{3/2}}$ (33)
$\frac{\partial\dot{f}}{\partial e}=\sqrt{\frac{M}{a^{3}}}\Big{(}\frac{2(1+e\cos{f})\cos{f}}{(1-e^{2})^{3/2}}+\frac{3e(1+e\cos{f})^{2}}{(1-e^{2})^{5/2}}\Big{)}$ (34)
$\frac{\partial\dot{f}}{\partial f}=-\frac{2\sqrt{M/a^{3}}(1+e\cos{f})e\sin{f}}{(1-e^{2})^{3/2}}$ (35)
$\frac{\partial\dot{r}}{\partial a}$ $=\Big{(}\frac{-2e}{1+e\cos{f}}-\frac{(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}\Big{)}\dot{e}$ (36)
@@
$\frac{\partial\dot{r}}{\partial e}$ $=\Big{(}\frac{-2e}{1+e\cos{f}}-\frac{(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}\Big{)}\dot{a}$ (37)
@@
$\frac{\partial\dot{r}}{\partial f}$ $=\frac{(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}\dot{a}$ (38)
@@
$+\Big{(}\frac{a(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}\Big{)}\frac{\partial\dot{f}}{\partial f}$ (39)
The terms found in equations 21 and 22 are given by
$\frac{\partial x}{\partial a}=\frac{\partial r}{\partial a}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial a}\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (40)
$\frac{\partial x}{\partial e}=\frac{\partial r}{\partial e}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial e}\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (41)
$\frac{\partial x}{\partial i}=r\sin{i}\sin{\Omega}\sin{(\omega+f)}$ (42)
$\frac{\partial x}{\partial\Omega}=-r\sin{\Omega}\cos{(\omega+f)}-r\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (43)
$\frac{\partial x}{\partial\omega}=-r\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\sin{\Omega}\cos{(\omega+f)}$ (44)
$\frac{\partial x}{\partial f}$ $=\frac{\partial r}{\partial f}\cos{\Omega}\cos{(\omega+f)}-r\cos{\Omega}\sin{(\omega+f)}$ (45)
@@
$\frac{\partial y}{\partial a}=\frac{\partial r}{\partial a}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial a}\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (46)
$\frac{\partial y}{\partial e}=\frac{\partial r}{\partial e}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial e}\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (47)
$\frac{\partial y}{\partial i}=-r\sin{i}\cos{\Omega}\sin{(\omega+f)}$ (48)
$\frac{\partial y}{\partial\Omega}=r\cos{\Omega}\cos{(\omega+f)}-r\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (49)
$\frac{\partial y}{\partial\omega}=-r\sin{\Omega}\sin{(\omega+f)}+r\cos{i}\cos{\Omega}\cos{(\omega+f)}$ (50)
$\frac{\partial y}{\partial f}$ $=\frac{\partial r}{\partial f}\sin{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\sin{(\omega+f)}$ (51)
@@
$\frac{\partial z}{\partial a}=\frac{\partial r}{\partial a}\sin{i}\sin{(\omega+f)}$ (52)
$\frac{\partial z}{\partial e}=\frac{\partial r}{\partial e}\sin{i}\sin{(\omega+f)}$ (53)
$\frac{\partial z}{\partial i}=r\cos{i}\sin{(\omega+f)}$ (54)
$\frac{\partial z}{\partial\Omega}=0.0$ (55)
$\frac{\partial z}{\partial\omega}=r\sin{i}\cos{(\omega+f)}$ (56)
$\frac{\partial z}{\partial f}=\frac{\partial r}{\partial f}\sin{i}\sin{(\omega+f)}+r\sin{i}\cos{(\omega+f)}$ (57)
$\frac{\partial\dot{x}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\cos{(\Omega)}\cos{(\omega+f)}-\frac{\partial r}{\partial a}\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (58)
@@
$\frac{\partial\dot{x}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial e}\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (59)
@@
$\frac{\partial\dot{x}}{\partial i}$ $=\dot{r}\sin{i}\sin{\Omega}\sin{(\omega+f)}+r\cos{i}\dot{I}\sin{\Omega}\sin{(\omega+f)}$ (60)
@@
$\frac{\partial\dot{x}}{\partial\Omega}$ $=-\dot{r}\sin{\Omega}\cos{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (61)
@@
$\frac{\partial\dot{x}}{\partial f}$ $=-\dot{r}\cos{\Omega}\sin{(\omega+f)}+r\sin{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (62)
@@
$\frac{\partial\dot{y}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial a}\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (63)
@@
$\frac{\partial\dot{y}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial e}\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (64)
@@
$\frac{\partial\dot{y}}{\partial i}$ $=-\dot{r}\sin{i}\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\dot{I}\cos{\Omega}\sin{(\omega+f)}$ (65)
@@
$\frac{\partial\dot{y}}{\partial\Omega}$ $=\dot{r}\cos{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (66)
@@
$\frac{\partial\dot{y}}{\partial\omega}$ $=-\dot{r}\sin{\Omega}\sin{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (67)
@@
$\frac{\partial\dot{y}}{\partial f}$ $=-\dot{r}\sin{\Omega}\sin{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (68)
@@
$\frac{\partial\dot{z}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\sin{i}\sin{(\omega+f)}+\frac{\partial r}{\partial a}\cos{i}\dot{I}\sin{(\omega+f)}$ (69)
@@
$\frac{\partial\dot{z}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\sin{i}\sin{(\omega+f)}+\frac{\partial r}{\partial e}\cos{i}\dot{I}\sin{(\omega+f)}$ (70)
@@
$\frac{\partial\dot{z}}{\partial i}$ $=\dot{r}\cos{i}\sin{(\omega+f)}-r\sin{i}\dot{I}\sin{(\omega+f)}$ (71)
@@
$\frac{\partial\dot{z}}{\partial\Omega}=0.0$ (72)
$\frac{\partial\dot{z}}{\partial\omega}$ $=\dot{r}\sin{i}\cos{(\omega+f)}+r\cos{i}\dot{I}\cos{(\omega+f)}$ (73)
@@
$\frac{\partial\dot{z}}{\partial f}$ $=\dot{r}\sin{i}\cos{(\omega+f)}+r\cos{i}\dot{I}\cos{(\omega+f)}$ (74)

### v2

@@ Appendix B Modifications to the equations of motion
Modifications to Mercury 6.2 in the way we are using them, were done by Dawson & Murray-Clay (2012) and Wolff et al. (2012) and are described in the Appendix of Wolff et al. (2012). We reiterate those modifications in this section with specifics and discuss additional modifications to the equations of motions to allow for artificial evolution of the phase angles: argument of pericenter and longitude of ascending node. Modifications to REBOUND are explained below as well. We note that these modifications are non-symplectic, even for a symplectic integrator, thus, the modifications must represent small perturbations on the planet for this to be valid. These modifications were tested for the MVS and HYBRID integrators within Mercury and for the WHFAST and MERCURIUS integrators within REBOUND; the tests were run for systems with a star and 1, 2, or 4 planets.
@@
    Then, we transform the coordinates back to democratic heliocentric with
    reb_integrator_mercurius_inertial_to_dh(r)
  • Adding the six new members to the structure REB_PAR TICLEVUSRX, VUSRY, VUSRZ, AUSRX, AUSRY, AUSRZ– caused the simulation archive to save simulation data incorrectly(e.g.,particle parameters were assigned to the wrong REB_PARTICLE member). To work around this, we save simulation outputs into txt or hdf5 files directly. Future modifications will provide a method to save the simulation with the simulation archive.
@@
B.1 Example of the Impact of a User-defined Velocity
Here we discuss how including both additional velocities and accelerations instead of just additional accelerations allows tighter control of the planets’ evolution. With both additional velocities and accelerations, the position and velocity of the planet are updated every timestep. Since a set of position and velocity vectors defines a set of six osculating orbital elements, the evolution is done in a consistent way and the elements can be evolved independently from each other. If we only update the 3 components of the velocity vector every timestep (i.e only additional acceleration present), then this does not map consistently to a change in 6 orbital elements. To build intuition for this, we provide an example of one planet orbiting a star. Figure 10 shows the orbital evolution of one planet whose eccentricity is forced to change (dashed line top middle panel) for two cases: (1) both additional velocities and additional accelerations are used (orange) and (2) only additional accelerations (blue). The planet’s eccentricity is exponentially damped (Equation 6c) with $\tau_{e}=1$ Myr and $\Delta e=0.2$. For case (1), the planet’s eccentricity evolves exactly as prescribed and the other orbital elements are not disturbed. For case (2), the planet’s eccentricity damps further than prescribed, and the semi-major axis also damps. This happens because the additional acceleration applied to change the eccentricity also introduces a torque that changes the energy and hence the semi-major axis.
Figure 10: Two examples of one planet’s orbital evolution when the eccentricity is forced to damp exponentially (dashed maroon line). When both the user-defined additional velocities and accelerations are used (orange), the planet’s orbital elements evolve independently from each other and only the eccentricity damps as prescribed. When only the user-defined accelerations are used (blue), the planet’s semi-major axis also damps, and the eccentricity evolves lower than prescribed.
B.2 Updating Position and Velocity in REBOUND
The osculating orbital elements that are used for the user velocity and acceleration are calculated once each time the acceleration is calculated (or updated). The order of operations in the mercurius algorithm depends on various integrator-dependent factors (see Rein et al. 2019), but in a typical path, the acceleration is calculated twice in a timestep. This integrator generally follows a kick-drift-kick algorithm, where the momentum is updated for half the timestep, the position is updated for the full timestep, and finally the momentum is updated for the remaining half timestep (see Rein et al., 2019, for a thorough discussion). The kick or “interaction” step in mercurius, integrates the planets’ motion due to interactions between the planets, whereas the drift or “Kepler” step integrates the dominant Keplerian motion around the star. In the publicly available REBOUND code version 4, the interaction step only updates the velocity of the particles given their acceleration, which includes the user-defined acceleration. We modified this function to also update the position of the particles given the user-defined velocities; therefore, the particles experience an extra push in the kick step. This modification is non-symplectic within a symplectic integrator, which is allowed as long as the additional perturbations are small compared to the gravitational effects of the star and other planets. We note that our implemented procedure always updates position and velocity at the same time based on the same set of orbital elements. We tested our modifications for integrators WHFAST and MERCURIUS, thus position and velocity are updated simultaneously in time. If one is interested in following these modifications for other integrators, one should be aware if this is appropriate for that integrator.
B.3 Partial Derivatives in User-defined Acceleration and Velocity
Here we present the partial derivatives that are in the user-defined acceleration and velocity (Equation 5), which are the additional terms used to artificially evolve a planet’s orbital elements. The partial derivatives for the cartesian coordinates ($x$, $y$, $z$, $\dot{x}$, $\dot{y}$, $\dot{z}$) are with respect to the orbital elements, since each orbital element is independent from each other and is time-evolving (i.e. $\dot{a}$, $\dot{e}$, $\dot{I}$, $\dot{\omega}$, $\dot{\Omega}$, $\dot{f}$ may be non-zero). The orbital elements in these equations are derived each timestep from the positions and velocities using standard formulas for osculating orbital elements. The variable rr—the distance of the planet from the central body—is itself a function of orbital elements, $r=\frac{a(1-e^{2})}{1+e\cos{f}}$, so we show partial derivatives for $r$ and $\dot{r}$ as well.
$\frac{\partial r}{\partial a}=\frac{1-e^{2}}{1+e\cos{f}}=\frac{r}{a}$ (B1)
$\frac{\partial r}{\partial e}=-\frac{2ae}{1+e\cos{f}}-\frac{a(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}$ (B2)
$\frac{\partial r}{\partial f}=\frac{a(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}$ (B3)
$\frac{\partial r\dot{f}}{\partial a}=\frac{-r\dot{f}}{2a}=-\frac{1}{2}\sqrt{\frac{M}{a^{3}}}\frac{1+e\cos{f}}{\sqrt{1-e^{2}}}$ (B4)
$\frac{\partial r\dot{f}}{\partial e}=\frac{r\dot{f}(e+\cos{f})}{(1-e^{2})(1+e\cos{f})}=\sqrt{\frac{M}{a}}\frac{e+e\cos{f}}{(1-e^{2})^{3/2}}$ (B5)
$\frac{\partial r\dot{f}}{\partial f}=\frac{-a\sqrt{M/a^{3}}e\sin{f}}{\sqrt{1-e^{2}}}$ (B6)
$\frac{\partial\dot{f}}{\partial a}=-3/2\frac{\sqrt{M/a^{5}}(1+e\cos{f})^{2}}{(1-e^{2})^{3/2}}$ (B7)
$\frac{\partial\dot{f}}{\partial e}=\sqrt{\frac{M}{a^{3}}}\Big(\frac{2(1+e\cos{f})\cos{f}}{(1-e^{2})^{3/2}}+\frac{3e(1+e\cos{f})^{2}}{(1-e^{2})^{5/2}}\Big)$ (B8)
$\frac{\partial\dot{f}}{\partial f}=-\frac{2\sqrt{M/a^{3}}(1+e\cos{f})e\sin{f}}{(1-e^{2})^{3/2}}$ (B9)
$\frac{\partial\dot{r}}{\partial a}$ $=\Big(\frac{-2e}{1+e\cos{f}}-\frac{(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}\Big)\dot{e}$ (B10)
@@ B.3 Partial Derivatives in User-defined Acceleration and Velocity
$\frac{\partial\dot{r}}{\partial e}$ $=\Big(\frac{-2e}{1+e\cos{f}}-\frac{(1-e^{2})\cos{f}}{(1+e\cos{f})^{2}}\Big)\dot{a}$ (B11)
@@
$\frac{\partial\dot{r}}{\partial f}$ $=\frac{(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}\dot{a}$ (B12)
@@
$+\Big(\frac{a(1-e^{2})e\sin{f}}{(1+e\cos{f})^{2}}\Big)\frac{\partial\dot{f}}{\partial f}$ (B13)
$\frac{\partial x}{\partial a}=\frac{\partial r}{\partial a}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial a}\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (B14)
$\frac{\partial x}{\partial e}=\frac{\partial r}{\partial e}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial e}\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (B15)
$\frac{\partial x}{\partial i}=r\sin{i}\sin{\Omega}\sin{(\omega+f)}$ (B16)
$\frac{\partial x}{\partial\Omega}=-r\sin{\Omega}\cos{(\omega+f)}-r\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (B17)
$\frac{\partial x}{\partial\omega}=-r\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\sin{\Omega}\cos{(\omega+f)}$ (B18)
$\frac{\partial x}{\partial f}$ $=\frac{\partial r}{\partial f}\cos{\Omega}\cos{(\omega+f)}-r\cos{\Omega}\sin{(\omega+f)}$ (B19)
@@
$\frac{\partial y}{\partial a}=\frac{\partial r}{\partial a}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial a}\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (B20)
$\frac{\partial y}{\partial e}=\frac{\partial r}{\partial e}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial e}\cos{i}\cos{\Omega}\sin{(\omega+f)}$ (B21)
$\frac{\partial y}{\partial i}=-r\sin{i}\cos{\Omega}\sin{(\omega+f)}$ (B22)
$\frac{\partial y}{\partial\Omega}=r\cos{\Omega}\cos{(\omega+f)}-r\cos{i}\sin{\Omega}\sin{(\omega+f)}$ (B23)
$\frac{\partial y}{\partial\omega}=-r\sin{\Omega}\sin{(\omega+f)}+r\cos{i}\cos{\Omega}\cos{(\omega+f)}$ (B24)
$\frac{\partial y}{\partial f}$ $=\frac{\partial r}{\partial f}\sin{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\sin{(\omega+f)}$ (B25)
@@
$\frac{\partial z}{\partial a}=\frac{\partial r}{\partial a}\sin{i}\sin{(\omega+f)}$ (B26)
$\frac{\partial z}{\partial e}=\frac{\partial r}{\partial e}\sin{i}\sin{(\omega+f)}$ (B27)
$\frac{\partial z}{\partial i}=r\cos{i}\sin{(\omega+f)}$ (B28)
$\frac{\partial z}{\partial\Omega}=0.0$ (B29)
$\frac{\partial z}{\partial\omega}=r\sin{i}\cos{(\omega+f)}$ (B30)
$\frac{\partial z}{\partial f}=\frac{\partial r}{\partial f}\sin{i}\sin{(\omega+f)}+r\sin{i}\cos{(\omega+f)}$ (B31)
$\frac{\partial\dot{x}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\cos{(\Omega)}\cos{(\omega+f)}-\frac{\partial r}{\partial a}\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B32)
@@
$\frac{\partial\dot{x}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\cos{\Omega}\cos{(\omega+f)}-\frac{\partial r}{\partial e}\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B33)
@@
$\frac{\partial\dot{x}}{\partial i}$ $=\dot{r}\sin{i}\sin{\Omega}\sin{(\omega+f)}+r\cos{i}\dot{I}\sin{\Omega}\sin{(\omega+f)}$ (B34)
@@
$\frac{\partial\dot{x}}{\partial\Omega}$ $=-\dot{r}\sin{\Omega}\cos{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B35)
@@
$\frac{\partial\dot{x}}{\partial f}$ $=-\dot{r}\cos{\Omega}\sin{(\omega+f)}+r\sin{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (B36)
@@
$\frac{\partial\dot{y}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial a}\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B37)
@@
$\frac{\partial\dot{y}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\sin{\Omega}\cos{(\omega+f)}+\frac{\partial r}{\partial e}\cos{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B38)
@@
$\frac{\partial\dot{y}}{\partial i}$ $=-\dot{r}\sin{i}\cos{\Omega}\sin{(\omega+f)}-r\cos{i}\dot{I}\cos{\Omega}\sin{(\omega+f)}$ (B39)
@@
$\frac{\partial\dot{y}}{\partial\Omega}$ $=\dot{r}\cos{\Omega}\cos{(\omega+f)}-r\sin{\Omega}\dot{\Omega}\cos{(\omega+f)}$ (B40)
@@
$\frac{\partial\dot{y}}{\partial\omega}$ $=-\dot{r}\sin{\Omega}\sin{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (B41)
@@
$\frac{\partial\dot{y}}{\partial f}$ $=-\dot{r}\sin{\Omega}\sin{(\omega+f)}-r\cos{\Omega}\dot{\Omega}\sin{(\omega+f)}$ (B42)
@@
$\frac{\partial\dot{z}}{\partial a}$ $=\frac{\partial\dot{r}}{\partial a}\sin{i}\sin{(\omega+f)}+\frac{\partial r}{\partial a}\cos{i}\dot{I}\sin{(\omega+f)}$ (B43)
@@
$\frac{\partial\dot{z}}{\partial e}$ $=\frac{\partial\dot{r}}{\partial e}\sin{i}\sin{(\omega+f)}+\frac{\partial r}{\partial e}\cos{i}\dot{I}\sin{(\omega+f)}$ (B44)
@@
$\frac{\partial\dot{z}}{\partial i}$ $=\dot{r}\cos{i}\sin{(\omega+f)}-r\sin{i}\dot{I}\sin{(\omega+f)}$ (B45)
@@
$\frac{\partial\dot{z}}{\partial\Omega}=0.0$ (B46)
$\frac{\partial\dot{z}}{\partial\omega}$ $=\dot{r}\sin{i}\cos{(\omega+f)}+r\cos{i}\dot{I}\cos{(\omega+f)}$ (B47)
@@
$\frac{\partial\dot{z}}{\partial f}$ $=\dot{r}\sin{i}\cos{(\omega+f)}+r\cos{i}\dot{I}\cos{(\omega+f)}$ (B48)